OpenStudy (lohy725):

Five added to 3 times a number is equal to 8 less than 5 times the number

OpenStudy (jonathan34):

ok

OpenStudy (jonathan34):

If the sum of a number and two is doubled, the result is six less than three times the number. Find the number. 2(n + 2) = 3n - 6 2n + 4 = 3n -6 n = 10

OpenStudy (jonathan34):

best reponse PLEASE

OpenStudy (jonathan34):

thanks

OpenStudy (lohy725):

The answers are: 3x + 5 = 5x – 8 3x – 4 = 5x + 8 4x + 3 = 8x – 5 8x + 5 = 4x + 3

OpenStudy (jonathan34):

10 divided by 2 ?

OpenStudy (lohy725):

I think i figured it out.. Thank you!

OpenStudy (jonathan34):

np